### v1

@@ IV-C External validation
IV-D Latent space visualization with t-SNE
@@
TABLE II: Artifact Detection Performance on the KidsBrainIT dataset. Best performance for each signal is highlighted in bold
(a) VAE-IF
(b) LSTM-ATTn
(c) XGBoost
Signal Metric Mean Std
BPm Sensitivity 0.802 0.219
Specificity 0.921 0.041
ICPm Sensitivity 0.700 0.388
Specificity 0.926 0.052
HRT Sensitivity 0.549 0.425
Specificity 0.855 0.227
Signal Metric Mean Std
BPm Sensitivity 0.767 0.175
Specificity 0.991 0.010
ICPm Sensitivity 0.695 0.361
Specificity 0.993 0.007
HRT Sensitivity 0.592 0.400
Specificity 0.916 0.243
Signal Metric Mean Std
BPm Sensitivity 0.774 0.163
Specificity 0.995 0.007
ICPm Sensitivity 0.672 0.355
Specificity 0.990 0.015
HRT Sensitivity 0.465 0.414
Specificity 0.915 0.210
(b) LSTM-ATTn
(c) XGBoost
IV-E Ablation study
We first evaluated the importance of feature extraction using the VAE. For this, we compared the performance of our model with a standard Isolation Forest model trained directly on the patient recordings. Before training, we transformed each sample into a feature vector containing the current sample and 14 previous samples, forming a feature vector of size 15 (i.e., $W=15$ as was used to train the VAE). Next, we analyzed the significance of the self-attention component in the encoder of the VAE. To do so, we removed the self-attention component and trained the model with the same set of hyperparameters estimated by BO. The results of these ablation studies are presented in Table III.
TABLE III: Performance Comparison of iForest and VAE-IF without attention, showing that both models achieve low performance compared with the VAE-IF with attention approach.
iForest VAE-IF (no attention)
Signal Metric Mean Std Mean Std
BPm Sensitivity 0.294 0.221 0.602 0.227
Specificity 0.973 0.020 0.920 0.040
ICPm Sensitivity 0.274 0.321 0.545 0.358
Specificity 0.959 0.030 0.937 0.042
HRT Sensitivity 0.270 0.366 0.395 0.391
Specificity 0.885 0.255 0.863 0.206
V Discussion
In this study, we proposed an unsupervised approach leveraging the power of VAE and iIForest for the detection of artifacts in physiological time series collected at minute-by-minute resolution in ICU settings. To assess its effectiveness, we conducted a comparative analysis against LSTM-Attention (LSTM-ATTn) and XGBoost, both of which are supervised models. Our results demonstrate that our proposed VAE-IF method achieves sensitivity levels comparable to LSTM-ATTn and XGBoost while maintaining high specificity. Notably, it can operate in circumstances where no labels about artifacts are available. Our results clearly demonstrate the ability of VAE-IF to accurately detect artifacts. Furthermore, we visualize the learned latent space representation using t-SNE, revealing that a VAE trained on raw data containing both clean and artifactual samples effectively disentangles clean from noisy instances. These findings highlight the potential of unsupervised methods for artifact detection in physiological time series data, offering a promising solution to the scarcity of labeled data.
@@ V Discussion
VI Limitations and future work
@@
VII Conclusion
We have successfully developed an unsupervised approach for artifact detection in minute-by-minute physiological recordings, utilizing real-world, routinely collected ICU data. Our approach has demonstrated performance on par with fully supervised methods, highlighting its efficacy and potential for practical implementation as it does not require any data labelling. Furthermore, we have validated the generalizability of our approach by testing it on sample patients from an external dataset, and the model predictions align closely with visual observations. Future studies involving diverse datasets will provide valuable insights and ensure the robustness of our approach in various clinical settings. Overall, our findings present a promising direction for advancing artifact detection methodologies in physiological time series data, with potential applications in clinical research.
Acknowledgment

### v2

@@ 4.3 External validation
4.4 Latent space visualization with t-SNE
@@
Table 3: Artifact Detection Performance on the KidsBrainIT dataset. Best performance for each signal is highlighted in bold.
Signal Metric VAE-IF LSTM-ATTn XGBoost ARIMA IF VAE-IF (no attention)
BPm Sens. 0.802(0.219) 0.767(0.175) 0.774(0.163) 0.606(0.238) 0.294(0.221) 0.602(0.227)
Spec. 0.921(0.041) 0.991(0.010) 0.995(0.007) 0.994(0.006) 0.973(0.020) 0.920(0.040)
ICPm Sens. 0.700(0.388) 0.695(0.361) 0.672(0.355) 0.323(0.316) 0.274(0.321) 0.545(0.358)
Spec. 0.926(0.052) 0.993(0.007) 0.990(0.015) 0.991(0.006) 0.959(0.030) 0.937(0.042)
HRT Sens. 0.549(0.425) 0.592(0.400) 0.465(0.414) 0.434(0.445) 0.270(0.366) 0.395(0.391)
Spec. 0.855(0.227) 0.916(0.243) 0.915(0.210) 0.922(0.231) 0.885(0.255) 0.863(0.206)
4.5 Ablation study
We first evaluated the importance of feature extraction using the VAE. For this, we compared the performance of our model with a standard Isolation Forest model trained directly on the patient recordings. Before training, we transformed each sample into a feature vector containing the current sample and 14 previous samples, forming a feature vector of size 15 (i.e., $W=15$ as was used to train the VAE). Next, we analyzed the significance of the self-attention component in the encoder of the VAE. To do so, we removed the self-attention component and trained the model with the same set of hyperparameters estimated by BO. The results presented in Table 3, show that both models achieve low performance compared with the VAE-IF with attention approach.
5 Discussion
In this study, we proposed a fully unsupervised hybrid method leveraging the power of VAE and IF for the detection of artifacts in physiological time series routinely collected at minute-by-minute resolution in ICU settings. To assess its effectiveness, we conducted a comparative analysis against LSTM-Attention (LSTM-ATTn) and XGBoost, both of which are supervised machine learnig models, and with a statistical approach using ARIMA. Our results demonstrate that our proposed VAE-IF method achieves sensitivity levels comparable to LSTM-ATTn and XGBoost while maintaining high specificity, with all three models significantly outperforming the standard ARIMA model used for comparison. Notably, our method can operate in circumstances where no labels about artifacts are available. Our results clearly demonstrate the ability of VAE-IF to accurately detect artifacts. Furthermore, we visualize the learned latent space representation using t-SNE, revealing that a VAE trained on raw data containing both clean and artifactual samples effectively disentangles clean from noisy instances. These findings highlight the potential of unsupervised methods for artifact detection in physiological time series data, offering a promising solution to the scarcity of labeled data.
@@ 5 Discussion
5.1 Limitations and future work
@@
6 Conclusion
We have successfully developed an unsupervised approach for artifact detection in minute-by-minute physiological recordings, utilizing real-world, routinely collected ICU data. Our approach has demonstrated performance on par with fully supervised methods, highlighting its efficacy and potential for practical implementation as it does not require any data labelling. Furthermore, we have validated the generalizability of our approach by testing it on sample patients from an external dataset, and the model predictions align closely with visual observations. Future studies involving diverse datasets will provide valuable insights and ensure the robustness of our approach to other clinical settings and the possibility to build on VAE-IF to develop an end-to-end method. Overall, our findings present a promising direction for advancing artifact detection methodologies in physiological time series data, with potential applications in clinical research.
Acknowledgements
@@
This research was funded in part by an EPSRC Doctoral Training Partnership PhD studentship to HH and The Carnegie Trust for the Universities of Scotland (RIG009251) to JE. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.